OpenStudy (ebonyserrano):

Three consecutive odd integers have a sum of 45. Find the integers. would that be 5,7, and 9?

OpenStudy (solomonzelman):

That would be incorrect, because 5+7+9=21 (not =45).

OpenStudy (solomonzelman):

Let your middle integer be \(\color{black}{\displaystyle x }\) (this is the second consecutive even integer you are looking for).

OpenStudy (solomonzelman):

Then, the smallest (the first) integer is \(\color{black}{\displaystyle x-2 }\) and the largest (the third) integer is \(\color{black}{\displaystyle x+2 }\).

OpenStudy (solomonzelman):

You know that these have to add to 45, so we set: \(\color{black}{\displaystyle (x-2)+x+(x+2)=45 }\) where \(\color{black}{\displaystyle x }\) denotes the middle (the 2nd) of the consecutive odd integers.

OpenStudy (solomonzelman):

Solve for \(x\), and from there you will know what the first odd integer and the third odd integer are.

OpenStudy (solomonzelman):

Example: Three consecutive odd integers have a sum of 39. Find the integers. We can call the middle integer \(x\). (With this definition, \((x-2)\) and \((x+2)\) are the first and third integers respectively.) Knowing that the sum of these three is equal to 36, we set the following equation. \(\color{black}{\displaystyle (x-2)+x+(x+2)=39}\) \(\color{black}{\displaystyle 3x=39\quad\Longrightarrow\quad x=13.}\) So, the middle number is \(13\), and therefore the remaining two are \(11\) and \(15\). (There is the answer to the example: \(11\), \(13\), \(15\).)